OpenStudy (anonymous):

Express in logarithmic form for the base. 4^2=16

Parth (parthkohli):

\( \color{Black}{\Rightarrow a^b = c \Longrightarrow \log_ac = b}\)

Parth (parthkohli):

Use this example.

Parth (parthkohli):

Here: \(4 = a\) \(2 = b\) \(16 = c\)

OpenStudy (anonymous):

alright i got log(4) 16=2

Parth (parthkohli):

Yes.

OpenStudy (maheshmeghwal9):

yeah of course u r right 100% So as parth said.\[\huge{4^2}=16.\]\[\huge{\color{blue}\log_4\color{green}{16}=\color{red}2}\]

OpenStudy (anonymous):

what about with this express in exponential form log(3)9=2

OpenStudy (maheshmeghwal9):

just reverse the process:)

OpenStudy (anonymous):

could you put it into an equation like above so i can work from it

Parth (parthkohli):

\( \color{Black}{\Rightarrow \log _a b = c \Longrightarrow a^c = b}\)

OpenStudy (maheshmeghwal9):